# UltraFusion: Ultra High Dynamic Imaging using Exposure Fusion

### v2

@@ Abstract
Capturing high dynamic range (HDR) scenes is one of the most important issues in camera design. Majority of cameras use exposure fusion technique, which fuses images captured by different exposure levels, to increase dynamic range. However, this approach can only handle images with limited exposure difference, normally 3-4 stops. When applying to very high dynamic scenes where a large exposure difference is required, this approach often fails due to incorrect alignment or inconsistent lighting between inputs, or tone mapping artifacts. In this work, we propose UltraFusion, the first exposure fusion technique that can merge input with 9 stops differences. The key idea is that we model the exposure fusion as a guided inpainting problem, where the under-exposed image is used as a guidance to fill the missing information of over-exposed highlight in the over-exposed region. Using under-exposed image as a soft guidance, instead of a hard constrain, our model is robust to potential alignment issue or lighting variations. Moreover, utilizing the image prior of the generative model, our model also generates natural tone mapping, even for very high-dynamic range scene. Our approach outperforms HDR-Transformer on latest HDR benchmarks. Moreover, to test its performance in ultra high dynamic range scene, we capture a new real-world exposure fusion benchmark, UltraFusion dataset, with exposure difference up to 9 stops, and experiments show that UltraFusion can generate beautiful and high-quality fusion results under various scenarios. An online demo is provided at project page.
Figure 1: Comparing our guided inpainting HDR imaging method with a state-of-the-art HDR reconstruction [26] and multi-exposure fusion [49] methods. Both scenes are selected from our captured new real-world benchmark. Left: night cityscape with large exposure difference. Right: afternoon street with motion-induced occlusion. Previous methods struggle to handle these scenes. By modeling HDR as inpainting problem, our method can produce visually appealing results without ghosting artifacts in these challenging scenes.
@@ 1 Introduction
High dynamic range (HDR) imaging is one of the fundamental problems in the modern camera design. Due to the hardware limitation, camera sensors have a much smaller dynamic range compared with real-world. To increase it, majority of HDR solutions merge multiple images, either same [9] or different exposure levels [15, 50, 56, 57, 35, 3, 25, 41, 44, 17]. Despite all recent advances, majority HDR algorithms can only bring limited increase in dynamic range. For example, HDR+ [9], the first HDR algorithm used by commercial cameras, can only robustly increase the dynamic range by 8 times (3 stops). Therefore, in this work, we study the following question: can we drastically increase the dynamic range of a camera by fusing two images with very large exposure difference, like 9 stops in Fig. 1?
@@
Figure 2: Visual comparison on directly utilizing ControlNet [64] and our UltraFusion. ControlNet struggles to fix a frame for reference, while our method fixes the over-exposed image as the reference and the under-exposed image as the guidance for inpainting, thereby avoiding artifacts.
@@
There are three advantages of this approach when handling large exposure difference. First, it follows the exposure fusion [31] setup. Unlike HDR fusion techniques, which first generate HDR image and then compress to a low-dynamic-range (LDR) image through tone mapping, exposure fusion directly generates LDR output, avoiding cascading errors. Second, the under-exposed image is used as a soft guidance, instead of a hard constraint. Therefore, UltraFusion is robust to alignment error (see the right scene of Fig. 1) and lighting variation (see the left scene of  Fig. 1). Third, image prior of the generative model ensures the natural look of the output image, reducing the potential artifacts.
To train a guided inpainting, a simple solution is to train a ControlNet [64] using two pre-aligned input images. However, such paradigm cannot handle dynamic HDR scenes, as ControlNet may not know which frame to choose as the reference, increasing the difficulty to fuse the information from over and under-exposed images. For instance, as shown in Fig. 2, ControlNet selects the over-exposed image as the reference frame in red boxed region, but chooses the under-exposed image as the reference frame in green boxed region, leading to substantial artifacts in the result. Additionally, as a generative model, it may inevitably generate fake image content, as indicated by green box in Fig. 2.
To address these challenges, the proposed UltraFusion is designed as follows. First, we warp the short-exposed image to the long-exposed one, and mask out the occluded regions. Then, we utilize the diffusion prior to inpaint the long-exposed image guided by partial short-exposed information. To make the network retain more details to generate guidance information, we propose a new decompose-and-fuse control branch, which eliminates the luminance component of the short-exposed image, extracting structure and color information instead, and employs a multi-scale cross-attention to improve the feature fusion with long-exposed image.
To evaluate the effectiveness of our framework, we capture 100 under/over-exposed image pairs, covering daytime, nighttime, indoor, outdoor scenes with local and global motion patterns. Experiments on both latest HDR imaging datasets and our captured benchmark demonstrate that, comparing to existing methods, our UltraFusion is more robust to scenes with large exposure differences and large motion, as shown in Fig. 1.
@@ 2.1 HDR imaging
Figure 3: The whole backbone of UltraFusion. Our method is a 2-stage framework, consisting of (a) pre-alignment stage and (b) guided inpainting stage. The first stage pre-aligns the under-exposed image $I_{ue}$ to the over-exposed image $I_{oe}$ and masks the occluded regions. In the subsequent guided inpainting stage, we propose a new decompose-and-fuse control branch to utilize the diffusion priors.
@@ 3.2 Guided inpainting stage
Different from general diffusion-based image enhancement, we propose a novel decompose-and-fuse control branch to inject two input images and diffusion latent as a control signal, as Fig. 10 (b) shows that naively injecting this information may not be able to guide diffusion to faithfully inpaint missing highlights obtained from the under-exposed image.
@@
Therefore, we decompose the underexposed image into the color and structure information, both of which are robust to brightness changes. Specifically, we use the normalized image as the structure component, similar to SSIM [47], as:
@@
where $Y_{ue}$ represents the luminance channel of $I_{ue}$ under YUV space, $\mu(\cdot)$ and $\sigma(\cdot)$ denote the mean intensity and standard deviation, respectively. The chroma channels (UV) are used as color information. The extracted structure and color information are further encoded using trained color and structure extractors (gray GE block in Fig. 4).
@@ 4 Experiment
Figure 5: (a) The data distribution of our benchmark. (b) The user study result on our benchmark.
@@
4.1 Experimental setting
Datasets. We evaluate our method on both static datasets and dynamic datasets. For evaluation, we use the MEFB dataset [65] with 100 static under/over-exposed image pairs and RealHDRV [40], a dynamic HDR deghosting test set containing 50 scenes with varying motion patterns.
@@ 4.1 Experimental setting
Implementation details. We leverage the generative prior encapsulated in Stable Diffusion V2.1 [38]. The decompose-and-fuse control branch (DFCB) is trained for 140k iterations with batch size $bs=32$ on 8 NVIDIA RTX 4090 GPUs. Adam is adopted as the optimizor and the learning rate is fixed to 0.0001. To adapt HDR reconstruction methods to 2 differently exposed inputs, we re-implement them by following their default settings.
Evaluation metrics. We utilize three widely-used non-reference metric MUSIQ [16], PAQ2PIQ [61] and HyperIQA [42] for quantitative comparison. Moreover, For static dataset [65], as no ground truths are provided, we select the task-specific MEF-SSIM [28] for structure retention evaluation. For dynamic dataset [40] with HDR ground truths, TMQI [60] is used to evaluate the performance of fidelity and naturalness. In addition, we conduct a user study on our Ultrafusion benchmark to perform subjective evaluation.
Figure 8: Visual comparisons on our captured UltraFusion benchmark.
Figure 9: Visual results on dynamic RealHDRV dataset [40].
@@ 4.2 Comparisons with HDR imaging methods
Evaluation on static dataset. We evaluate the fusion performance on the MEFB dataset [65], focusing on large exposure differences. As shown in Fig. 6, our method outperforms other methods on all three non-reference metrics (MUSIQ, PAQ2PIQ, and HyperIQA). Specifically, our proposed UltraFusion achieves 2.06 gain in terms of MUSIQ compared to HSDS-MEF. In terms of MEF-SSIM, as shown in  Fig. 6, our baseline model (ControlNet [64]) achieves better image quality but lacks fidelity, while HSDS-MEF [49] retains more information from inputs at the cost of quality. Our UltraFusion outperforms most algorithms, and achieves similar fidelity scores compared to HSDS-MEF and TC-MoA but with much higher image quality (non-reference metrics), indicating the best trade-off between visual quality and information preservation. The qualitative comparison in Fig. 7 further validates this. In contrast, HDR reconstruction methods (HDR-Transformer and SCTNet) often miss some detail in highlights and MEF methods (HSDS-MEF and TC-MoA) introduce unnatural transition from brighter to dark regions.
@@
Evaluation on our UltraFusion benchmark. At last, the evaluation on our benchmark validates the robustness of UltraFusion in the most challenging scenes. On all four non-reference metrics, our method outperforms compititors by a large margin, as shown in Tab. 2. The qualitative comparison in Fig. 8 is also consistent with the quantitative metrics.
For example, in the red zoom-in patch, integrating the highly bright sun from the under-exposed image into the over-exposed image is extremely challenging. Other methods fail to maintain the shape of the sun or preserve high contrast in the fused region, while our method naturally reconstructs the sun, preserving its appearance and visual-pleasing tone of the whole image. Moreover, we conduct a user study on our proposed benchmark. Specifically, we randomly select 20 scenes from our benchmark and invite 136 different users to participate. For each scene, each user is asked to compare our method with a randomly chosen baseline. The user study in Fig. 5 (b) indicates our method is more favored by users than competitors. This outcome aligns with the non-reference metric evaluation, showing that our method produces more natural images.
@@ 4.3 Ablation studies
To validate the effectiveness of our UltraFusion, we conduct ablation studies on key components, followed by an in-depth exploration of the designs of decompose-and-fuse control branch.
Key components. First, we perform the ablation study of the proposed key components, including the alignment strategy and decompose-and-fuse control branch, on the RealHDRV [40] dataset, as shown in Tab. 3. We first remove our alignment strategy, opting instead to train the model directly on the original dataset and input differently exposed image pairs without coarse alignment. Without the pre-alignment stage and the data synthesis pipeline designed for it, performance drops significantly in terms of TMQI. Fig. 10 (a) also shows that the model fails to implicitly align large motions. Then, we replace the decompose-and-fuse control branch (DFCB) with the vanilla ControlNet [64], and it fails to fuse features with large exposure differences, which leads to the loss of details (see  Fig. 10 (b)) and a decrease in TMQI.
@@
Decompose-and-fuse control branch. We also evaluate how the form of soft guidance provided by the under-exposed image impacts the recovery of the highlight regions. When using the RGB under-exposed image $I_{ue}$ as guidance, the model will ignore some details in the output (Fig. 11 (c)). Replacing $I_{ue}$ with its structure information $S_{ue}$ retains more details (Fig. 11 (d)), but without generating vivid color. By incorporating both under-exposed color information $C_{ue}$ and structure information $S_{ue}$, the reconstructed results can maintain more details and color consistency (see Fig. 11 (e)). At last, multi-scale cross-attention further improves the fusion result (see Fig. 11 (f)).

### v3

@@ Abstract
Capturing high dynamic range (HDR) scenes is one of the most important issues in camera design. Majority of cameras use exposure fusion, which fuses images captured by different exposure levels, to increase dynamic range. However, this approach can only handle images with limited exposure difference, normally 3-4 stops. When applying to very high dynamic range scenes where a large exposure difference is required, this approach often fails due to incorrect alignment or inconsistent lighting between inputs, or tone mapping artifacts. In this work, we propose UltraFusion, the first exposure fusion technique that can merge inputs with 9 stops differences. The key idea is that we model exposure fusion as a guided inpainting problem, where the under-exposed image is used as a guidance to fill the missing information of over-exposed highlights in the over-exposed region. Using an under-exposed image as a soft guidance, instead of a hard constraint, our model is robust to potential alignment issue or lighting variations. Moreover, by utilizing the image prior of the generative model, our model also generates natural tone mapping, even for very high-dynamic range scenes. Our approach outperforms HDR-Transformer on latest HDR benchmarks. Moreover, to test its performance in ultra high dynamic range scenes, we capture a new real-world exposure fusion benchmark, UltraFusion dataset, with exposure differences up to 9 stops, and experiments show that UltraFusion can generate beautiful and high-quality fusion results under various scenarios. Code and data will be available at https://openimaginglab.github.io/UltraFusion.
Figure 1: Comparing our guided inpainting HDR imaging method with a state-of-the-art HDR reconstruction [25] and multi-exposure fusion [48] methods. Both scenes are selected from our captured new real-world benchmark. Left: night cityscape with large exposure difference. Right: afternoon street with motion-induced occlusion. Previous methods struggle to handle these scenes. By modeling HDR as an inpainting problem, our method can produce visually appealing results without ghosting artifacts in these challenging scenes.
@@ 1 Introduction
High dynamic range (HDR) imaging is one of the fundamental problems in the modern camera design. Due to hardware limitation, camera sensors have a much smaller dynamic range compared with the real-world. To increase it, majority of HDR solutions merge multiple images, either with the same [9] or different exposure levels [15, 49, 56, 57, 34, 3, 24, 40, 43, 17]. Despite all recent advances, majority HDR algorithms can only bring limited increase in dynamic range. For example, HDR+ [9], the first HDR algorithm used by commercial cameras, can only robustly increase the dynamic range by 8 times (3 stops). Therefore, in this work, we study the following question: can we drastically increase the dynamic range of a camera by fusing two images with very large exposure difference, like 9 stops in Fig. 1?
@@
Figure 2: Visual comparison on directly utilizing ControlNet [65] and our UltraFusion. Without pre-aligned data, ControlNet struggles to fix a frame for reference. However, our method fixes the over-exposed image as the reference and the under-exposed image as the guidance for inpainting, thereby avoiding artifacts.
@@
There are three advantages of this approach when handling large exposure difference. First, it follows the exposure fusion [30] setup. Unlike HDR fusion techniques, which first generate HDR image and then compress to a low-dynamic-range (LDR) image through tone mapping, exposure fusion directly generates LDR output, avoiding cascading errors. That being said, our method outputs a tone-mapped LDR image instead of a linear HDR. Second, the under-exposed image is used as a soft guidance, instead of a hard constraint. Therefore, UltraFusion is robust to alignment error (see the right scene of Fig. 1) and lighting variation (see the left scene of  Fig. 1). Third, image prior of the generative model ensures the natural look of the output image, reducing the potential artifacts.
To train a guided inpainting, a simple solution is to train a ControlNet [65] using two input images. However, such paradigm cannot handle dynamic HDR scenes, as ControlNet may not know which frame to choose as the reference, increasing the difficulty to fuse the information from over and short-exposed images. For instance, as shown in Fig. 2, ControlNet selects the over-exposed image as the reference frame in red boxed region, but chooses the under-exposed image as the reference frame in green boxed region, leading to substantial artifacts in the result. Additionally, as a generative model, it may inevitably generate fake image content, as indicated by green box in Fig. 2.
To address these challenges, we design UltraFusion as follows. First, we warp the short-exposed image to the long-exposed one, and mask out the occluded regions. Then, we utilize the diffusion prior to inpaint the long-exposed image guided by partial short-exposed information. To make the network retain more details to generate guidance information, we propose a new decompose-and-fuse control branch, which eliminates the luminance component of the short-exposed image, extracting structure and color information instead, and employs a multi-scale cross-attention to improve the feature fusion with long-exposed image. Second, as there is no existing large-scale training data for exposure fusion of dynamic scene, we also propose a novel training data synthesis pipeline, utilizing existing high-quality multi-exposure datasets (pre-aligned) of static scene and video datasets. At last, to ensure the generated output maintains the fidelity to the reality, we also train an additional fidelity control branch, using the same decompose-and-fuse strategy and multi-scale cross-attention.
At last, to evaluate the effectiveness of our framework, we capture 100 under/over-exposed image pairs, covering daytime, nighttime, indoor, outdoor scenes with local and global motion patterns. Experiments on both latest HDR imaging datasets and our captured benchmark demonstrate that, comparing to existing methods, our UltraFusion is more robust to scenes with large exposure differences and large motion, as shown in Fig. 1.
@@ 2.1 HDR imaging
Figure 3: The whole backbone of UltraFusion. Our method is a 2-stage framework, consisting of (a) pre-alignment stage and (b) guided inpainting stage. The first stage pre-aligns the under-exposed image $I_{ue}$ to the over-exposed image $I_{oe}$ and masks the occluded regions. In the subsequent guided inpainting stage, we propose a new decompose-and-fuse control branch to utilize the diffusion priors, together with a fidelity control branch for fidelity keeping.
@@ 3.2 Guided inpainting stage
The main differences between our solution and general diffusion-based image enhancement are two-fold. First, we propose a novel decompose-and-fuse control branch to inject two input images and diffusion latent as a control signal, as Fig. 11 (b) shows that naively injecting this information may not be able to guide diffusion to faithfully inpaint missing highlights obtained from the under-exposed image. Second, we train an additional fidelity control branch to provide faithful structure and color information for the decoding process via shortcuts. Details are described below.
@@
Therefore, we decompose the under-exposed image into the color and structure information, both of which are robust to brightness changes. Specifically, we use the normalized image as the structure component, similar to SSIM [46], as:
@@
where $Y_{ue}$ represents the luminance channel of $I_{ue}$ under YUV space, $\mu(\cdot)$ and $\sigma(\cdot)$ denote the mean intensity and standard deviation, respectively. The chroma channels (UV) are used as color information. The extracted structure and color information are further encoded using trained color and structure extractors (gray GE block in Fig. 4). Following [65], we implement GE using several simple convolution layers to extract multi-scale features.
@@
Fidelity control branch. Even with the control block, we sometimes observe undesirable modifications of texture introduced by VAE. An example is shown in Fig. 11 (c). Therefore, to further improve fidelity, we design a Fidelity Control Branch (FCB), inspired by [47]. FCB mitigates texture distortions by injecting features into the VAE decoder. It has a similar architecture as the decompose-and-fuse control branch, with two main differences: 1) the main extractor of FCB adopts the same structure as the VAE Encoder, rather than the denoising U-Net, to provide corresponding shortcuts to the VAE Decoder, with adjustments to the soft guidance extractor as well, and 2) the main extractor of the FCB directly takes the over-exposed image as input. To train the fidelity control branch, we freeze the pre-trained VAE Encoder and Decoder and encode the ground truth $I_{gt}$ to latent space, simulating the denoised latent $z_{0}$ during the inference. Then we input the corresponding over-exposed image and under-exposed image to the fidelity control branch to extract faithful features. At last, the VAE Decoder decodes the compressed latent to a reconstructed image $\hat{I}_{gt}$. We adopt the reconstruction loss $||\hat{I}_{gt}-I_{gt}||_{1}$ as an additional loss term.
Figure 5: The illustration of our training data synthesis pipeline.
3.3 Training data synthesis
Preparing the training data for the proposed guided inpainting network is challenging. To train the model, ideally we need a large-scale HDR dataset that 1) covers different dynamic scenes, 2) has large exposure variance up to 9 stops, and 3) contains ground truth fusion result. However, no existing dataset satisfy all requirements.
To solve this issue, we propose a novel training data synthesis pipeline. Specifically, as shown in Fig. 5, we first randomly sample an image sequence with $N$ frames from a video dataset [55]. To model large motion, we select the first and the last frames. Then, similar to the pre-alignment stage, we estimate the bidirectional optical flow between two selected frames using [42] and compute a pseudo occlusion mask via forward-backward consistency check. Subsequently, we randomly sample an under-exposed image patch from the high-quality static dataset [1] (ground truths are pre-aligned), resize the pseudo occlusion mask to match the patch size, and mask out the pseudo occluded region to synthesize a pseudo pre-aligned output. With our synthesized training data, our UltraFusion learns to handle dynamic scenes with only static multi-exposure image pairs.
@@ 4 Experiment
4.1 Experimental setting
Datasets. We utilize SICE [1] dataset and Vimeo-90K [55] dataset to synthesize our training data. Following previous works [1, 22], we select images with the highest and lowest brightness in each exposure bracket of the SICE dataset as inputs. We evaluate our method on both static datasets and dynamic datasets. For evaluation, we use the MEFB dataset [66] with 100 static under/over-exposed image pairs and RealHDRV [39], a dynamic HDR deghosting test set containing 50 scenes with varying motion patterns.
Figure 6: (a) The data distribution of our benchmark. Coordinate value of exposure time represents the upper boundary. (b) The user study result on our benchmark.
@@ 4.1 Experimental setting
Implementation details. We leverage the generative prior encapsulated in Stable Diffusion V2.1 [37]. The decompose-and-fuse control branch (DFCB) is trained for 140k iterations with batch size $bs=32$ on 8 NVIDIA RTX 4090 GPUs. We also train the fidelity control branch for 1000K iteration with batch size $bs=1$ on a single NVIDIA RTX 4090 GPU. Adam is adopted as the optimizor and the learning rate is fixed to 0.0001. To adapt HDR reconstruction methods to 2 differently exposed inputs, we re-implement them by following their default settings.
Evaluation metrics. We utilize four widely used non-reference metric MUSIQ [16], DeQA-Score [62], PAQ2PIQ [61] and HyperIQA [41] for quantitative comparison. Moreover, for static dataset [66], as no ground truths are provided, we select the task-specific MEF-SSIM [27] for structure retention evaluation. For dynamic dataset [39] with HDR ground truths, TMQI [60] is used to evaluate the performance of fidelity and naturalness. In addition, we conduct a user study on our Ultrafusion benchmark to perform subjective evaluation.
@@ 4.2 Comparisons with HDR imaging methods
Figure 9: Visual comparisons on our captured UltraFusion benchmark.
Evaluation on static dataset. We evaluate the fusion performance on the MEFB dataset [66], focusing on large exposure differences. As shown in Fig. 7, our method outperforms other methods on all four non-reference metrics (MUSIQ, DeQA-Score, PAQ2PIQ, and HyperIQA). Specifically, our proposed UltraFusion achieves 2.06 gain in terms of MUSIQ compared to HSDS-MEF. In terms of MEF-SSIM, as shown in  Fig. 7, our baseline model (ControlNet [65]) achieves better image quality but lacks fidelity, while HSDS-MEF [48] retains more information from inputs at the cost of quality. Our UltraFusion outperforms most algorithms, and achieves similar fidelity scores compared to HSDS-MEF and TC-MoA but with much higher image quality (non-reference metrics), indicating the best trade-off between visual quality and information preservation. The qualitative comparison in Fig. 8 further validates this. In contrast, HDR reconstruction methods (HDR-Transformer and SCTNet) often miss some detail in highlights and MEF methods (HSDS-MEF and TC-MoA) introduce unnatural transition from brighter to dark regions.
@@
Figure 10: Visual results on dynamic RealHDRV dataset [39].
Evaluation on our UltraFusion benchmark. At last, the evaluation on our benchmark validates the robustness of UltraFusion in the most challenging scenes. On all four non-reference metrics, our method outperforms compititors by a large margin, as shown in Tab. 2. The qualitative comparison in Fig. 9 is also consistent with the quantitative metrics. For example, in the red zoom-in patch, integrating the highly bright sun from the under-exposed image into the over-exposed image is extremely challenging. Other methods fail to maintain the shape of the sun or preserve high contrast in the fused region, while our method naturally reconstructs the sun, preserving its appearance and visual-pleasing tone of the whole image. Moreover, we conduct a user study on our proposed benchmark. Specifically, we randomly select 20 scenes from our benchmark and invite 136 different users to participate. For each scene, each user is asked to compare our method with a randomly chosen baseline. The user study in Fig. 6 (b) indicates our method is more favored by users than competitors. This outcome aligns with the non-reference metric evaluation, showing that our method produces more natural images. More visual comparisons are available in the supplementary.
@@ 4.3 Ablation studies
To validate the effectiveness of our UltraFusion, we conduct ablation studies on the three key components, followed by an in-depth exploration of the designs of them.
Key components. First, we perform the ablation study of the proposed three key components, including the alignment strategy, decompose-and-fuse control branch, and fidelity control branch, on the RealHDRV [39] dataset, as shown in Tab. 3. We first remove our alignment strategy, opting instead to train the model directly on the original SICE [1] dataset and input differently exposed image pairs without coarse alignment. Without the pre-alignment stage and the data synthesis pipeline designed for it, performance drops significantly in terms of TMQI. Fig. 11 (a) also shows that the model fails to implicitly align large motions. Then, we replace the decompose-and-fuse control branch (DFCB) with the vanilla ControlNet [65], and it fails to fuse features with large exposure differences, which leads to the loss of details (see  Fig. 11 (b)) and a decrease in TMQI. Finally, when we exclude the fidelity control branch (FCB), UltraFusion loses the capability to maintain the detailed structure and vivid color, as indicated by Fig. 11 (c).
@@
Figure 12: Effectiveness of our alignment strategy.
Alignment strategy. Second, we perform ablation of the proposed alignment strategy, which is the most critical design for our framework. Our alignment strategy mainly consists of a data synthesis pipeline at training time and a pre-alignment module at testing time. Without both of them, the whole model degrades to a multi-exposure fusion method as shown in Fig. 12 (c). Performing pre-alignment can reduce artifacts to some extent (see Fig. 12 (d)), but may still be prone to alignment errors. Our data synthesis pipeline can improves the robustness of our algorithm to unaligned conditions, but motion in the occluded regions still cannot be solved, as shown in Fig. 12. After combining data synthesis pipeline and pre-alignment stage, our algorithm demonstrates strong capability towards dynamic scenes.
Decompose-and-fuse control branch. At last, we evaluate how the form of soft guidance provided by the under-exposed image impacts the recovery of the highlight regions. When using the RGB under-exposed image $I_{ue}$ as guidance, the model will ignore some details in the output (Fig. 13 (c)). Replacing $I_{ue}$ with its structure information $S_{ue}$ retains more details (Fig. 13 (d)), but without generating vivid color. By incorporating both under-exposed color information $C_{ue}$ and structure information $S_{ue}$, the reconstructed results can maintain more details and color consistency (see Fig. 13 (e)). At last, multi-scale cross-attention further improves the fusion result (see Fig. 13 (f)).